OpenStudy (anonymous):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15 If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250.

OpenStudy (anonymous):

what kind of graph would i use?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Okay so we have \(\large 85 \le 15c + 85 \le 250\) 85 because this is the fixed cost of the bracelet with no charms 15c because this is the cost of each charm and they save anywhere between 85 and 250 so this would shade all the amounts of charms they can buy

OpenStudy (anonymous):

like i would use a bar gragh?

OpenStudy (johnweldon1993):

Quite frankly I would just do a number line... We need to simplify that a bit more... \[\large 85 \le 15c + 85 \le 250\] \[\large 0 \le 15c \le 165\] \[\large 0 \le c \le 11\] This shows that you can buy anywhere from 0 to 11 charms for this bracelet..